# CORE-BEHRT A Carefully Optimized and Rigorously Evaluated BEHRT

### v2

@@ 5.1 Evaluation Approach/Study Design
Before setting up our experiments, we investigated the variance in PT and FT by running the base model multiple times. We found that both yielded stable performances with standard deviations of $6.6\cdot 10^{-4}$ and $3.2\cdot 10^{-4}$ for PT validation loss and FT validation AUROC, respectively. We investigated the source of variance in FT by conducting a leave-2-out CV; see Appendix E for details. The analysis revealed that the mean difference within test sets is $7.9\cdot 10^{-4}$ and pairwise difference for the means across test sets is $4.1\cdot 10^{-3}\,$, showcasing that variation is mainly attributed to the test set. These results gave us sufficient confidence to justify a single PT, followed by a five-fold CV. To ensure robust evaluation, we intentionally used larger-than-usual test sets.
5.2 Optimization Experiments
@@
Technical Components
Masking Ratios
We found that an MR of 20% was optimal, showing higher AUROCs for each of the three tasks compared to the base MR of 15%, as seen in Appendix I. At the same time, we observed that changes in masking ratios did not result in substantially different average AUROCs, and in general, the impact was inconsistent across the downstream tasks.
Pooling Strategies
The largest improvement in pooling strategies comes from the BiGRU head, but we show in Appendix I that most techniques performed better than the standard CLS token. Even simple aggregation functions, such as mean and max pooling, which use no additional trainable parameters, showed general improvements compared to CLS.
5.3 Generalization Results
@@
5.4 Maximizing Performance
The impact of increasing the CV set on model performance varies by condition, as seen in the comparison of Table 2 and Table 3. For breast cancer and preterm birth, performance increases are notable, rising by 2.4% and 7.1% in AUROC, respectively. In contrast, the AUROCs for stroke and schizophrenia remain largely unchanged despite the significant increases in patient numbers (3.6 times for stroke and 2.5 times for schizophrenia). This lack of improvement suggests a performance saturation at lower patient counts, supported by their already high baseline AUROCs of 70% and 86%, respectively.
Table 3: Results obtained with CORE-BEHRT model for an increased cross-validation set on select conditions. AUROC is given with standard deviation computed from the five folds. The same test sets as described in Table 2 was used.
@@ 5.4 Maximizing Performance
Patients in CV [thousands] Prevalence [%] AUROC [%]
Breast Cancer ($60+$/f.) 124.3 5.2 63.9$\pm$0.3
Preterm Birth 100.0 13.7 63.7$\pm$0.2
Schizophrenia ($20-30$) 89.2 1.4 85.9$\pm$0.6
Stroke ($60+$) 276.2 1.5 70.7$\pm$0.5
6 Discussion
@@
Increasing the size of our FT sets to maximize performance showed limited improvements for certain tasks despite the large data increase. This suggests that performance may saturate beyond a certain number of patients, dependent on the task. This saturation might happen faster at higher base scores, as evidenced by the initial high performance for schizophrenia. This indicates that more patients might not always be the solution, but data representation and technical components might be a more promising path for these scenarios.
@@
We show that improved data representation significantly boosts performance, with technical components providing an additional improvement. However, the consistent effects of these modifications remain unexplained, highlighting an important area for future research to uncover and analyze the underlying patterns.
Appendix A Time2vec Implementation
@@
Appendix B P-value Computation
@@
Appendix C Data processing
@@ Appendix D Hyperparameters
Appendix E Leave-2-Out Cross Validation
@@
Appendix F Cohort Selection
@@
Appendix G Generalization Cohort Statistics
@@
Appendix H Age Distribution for Generalization Cohort
@@
Appendix I Results of Generalization Experiments
@@
Appendix J Results of Optimization Experiments
@@
Appendix K Definitions of Conditions

### v4

@@ E.1 Evaluation Approach/Study Design
Before setting up our experiments, we investigated the variance in PT and FT by running the base model multiple times. We found that both yielded stable performances with standard deviations of $6.6\cdot 10^{-4}$ and $3.2\cdot 10^{-4}$ for PT validation loss and FT validation AUROC, respectively. We investigated the source of variance in FT by conducting a leave-2-out CV; see Appendix K for details. The analysis revealed that the mean difference within test sets is $7.9\cdot 10^{-4}$ and pairwise difference for the means across test sets is $4.1\cdot 10^{-3}\,$, showcasing that variation is mainly attributed to the test set. These results gave us sufficient confidence to justify a single PT, followed by a five-fold CV. To ensure robust evaluation, we intentionally used larger-than-usual test sets.
E.2 Optimization Experiments
@@
Technical Components
Masking Ratios
We found that an MR of 20% was optimal, showing higher AUROCs for each of the three tasks compared to the base MR of 15%, as seen in Appendix Q. At the same time, we observed that changes in masking ratios did not result in substantially different average AUROCs, and in general, the impact was inconsistent across the downstream tasks.
Pooling Strategies
The largest improvement in pooling strategies comes from the BiGRU head, but we show in Appendix Q that most techniques performed better than the standard CLS token. Even simple aggregation functions, such as mean and max pooling, which use no additional trainable parameters, showed general improvements compared to CLS.
@@
E.3 Generalization Results
@@
The impact of increasing the CV set can be seen in Appendix L. Breast cancer and preterm birth has notable AUROC increases, while stroke and schizophrenia remain largely unchanged despite the significant increases in patient numbers. This suggests a performance saturation at lower patient counts, supported by their high baseline AUROCs.
E.4 Out-of-Time Evaluation
When mimicking a real-world setting by performing an out-of-time evaluation, the size of the train and test sets increases dramatically and the prevalence decreases. Most tasks see a substantial performance increase, while death sees a small decrease and stroke a larger decrease as detailed in Table 3. Overall, this showcases the model’s ability to handle temporal shifts well.
Table 3: Results obtained mimicking a real-world setting, performing out-of-time evaluation, with train data being censored at 01/2020 and the test set contains outcomes from 09/2021 onwards. For comparison, we provide the difference in mean AUROC to the generalization results from Table 2.
@@ E.4 Out-of-Time Evaluation
Patients train/test [thousands] Prevalence in test [%] AUROC [%] $\Delta$ AUROC [%]
Breast Cancer ($60+$/f.) 72.0/262.9 0.9 64.4 $+2.0$
Stroke ($60+$) 174.3/516.8 0.2 67.1 $-3.1$
Diabetes 593.9/1,657.1 0.9 85.6 $+0.8$
Death ($60+$) 281.3/441.0 9.3 87.4 $-0.3$
Schizophrenia ($20-30$) 59.0/219.9 0.1 87.7 $+1.7$
Appendix F Discussion
@@
Increasing the FT set size yielded limited improvements for some tasks, suggesting performance saturation beyond a certain patient number. This saturation occurs faster with higher base scores, as seen with schizophrenia. Thus, adding more patients may not always help; instead, focusing on data representation and technical components could be more effective.
The out-of-time evaluation shows the model’s robustness to temporal shifts, maintaining high scores even with unseen time periods, likely due to the time2vec embedding. We believe this method is appropriate, as it mirrors real-world scenarios where models are used without constant retraining. The performance decrease on stroke might be caused by the larger prediction window with a minimum of three months, while the generalization experiment was performed with a 14-day prediction window.
@@
Our study focuses only on BERT-based EHR models, ignoring the classical models like Linear Regression, Random Forest, and Gradient Boosted Trees (XGBoost). While BERT-based EHR models have proven to beat simpler models, they have yet to be tested vs. XGBoost. We simply focus on BERT-based models as we believe these models have the most potential due to their versatility and scalability, as evident in NLP and computer vision. These models also have the potential to be multimodal in the future, as seen with LLMs, making them a strong candidate model for the future.
We show that improved data representation significantly boosts performance, with technical components providing an additional improvement. Although we only tested some modifications, we hope they act as representative examples and provide valuable insights into their respective areas. The consistent effects of these modifications remain unexplained, highlighting an important area for future research to uncover and analyze the underlying patterns.
Appendix G Time2vec Implementation
@@
Appendix H P-value Computation
@@
Appendix I Data processing
@@
Appendix K Leave-2-Out Cross Validation
@@
Appendix L Maximizing Performance
Table 4: Results obtained with CORE-BEHRT model for an increased cross-validation set on select conditions. AUROC is given with standard deviation computed from the five folds. The same test sets as described in Table 2 were used.
Condition
(Age Group/Sex)
Patients in CV [thousands] Prevalence [%] AUROC [%]
Breast Cancer ($60+$/f.) 124.3 5.2 63.9$\pm$0.3
Preterm Birth 100.0 13.7 63.7$\pm$0.2
Schizophrenia ($20-30$) 89.2 1.4 85.9$\pm$0.6
Stroke ($60+$) 276.2 1.5 70.7$\pm$0.5
The impact of increasing the CV set on model performance varies by condition, as shown in Table 2 and Table 4. For breast cancer and preterm birth, AUROC increases by 2.4% and 7.1%, respectively. However, AUROCs for stroke and schizophrenia remain largely unchanged despite a 3.6-fold and 2.5-fold increase in patient numbers. This suggests performance saturation at lower patient counts, given their high baseline AUROCs of 70% and 86%.
Appendix M Out of Time Evaluation
Figure 3: This figure illustrates the out-of-time fine-tuning and evaluation process. Training used data only up to 01/2020. The remaining three years were split for validation and testing: validation labels covered outcomes from 03/2020 to 06/2021, and testing labels covered outcomes from 09/2021 onward. Patients with outcomes before these dates in their respective sets were excluded.
Appendix N Cohort Selection
@@
Appendix O Generalization Cohort Statistics
@@
Appendix P Age Distribution for Generalization Cohort
@@
Appendix Q Results of Generalization Experiments
@@
Appendix R Results of Optimization Experiments
@@
Appendix S Definitions of Conditions